# Fairy2i: Training Complex LLMs from Real LLMs with All Parameters in $\{\pm 1,\pm i\}$

### v2

@@ Abstract
Large language models (LLMs) have revolutionized artificial intelligence, yet their massive memory and computational demands necessitate aggressive quantization, increasingly pushing representations toward the theoretical limit of a single bit. While complex-valued LLMs, such as iFairy, offer a superior chance for low-bit representation compared to real-valued counterparts, they require training from scratch, preventing the utilization of the vast ecosystem of pre-trained real-valued foundation models. Here we present Fairy2i, a universal framework that transforms pre-trained real-valued layers into an equivalent widely-linear complex form, enabling extremely low-bit quantization while reusing existing checkpoints. By proving a lossless mathematical equivalence between real and widely-linear maps, we convert standard Transformers into the complex domain and employ a phase-aware quantization scheme with a highly efficient codebook of fourth roots of unity ($\{\pm 1,\pm i\}$). Furthermore, we introduce a recursive residual quantization mechanism that iteratively minimizes quantization error, allowing inference to proceed via efficient multiplication-free accumulation. We demonstrate that Fairy2i restores the performance of LLaMA-2 7B at an effective 2-bit precision to levels nearly comparable with full-precision baselines, significantly outperforming state-of-the-art real-valued binary and ternary quantization methods. This work bridges the gap between the representational efficiency of complex-valued arithmetic and the practical utility of pre-trained models, paving a new way for efficient inference on commodity hardware.
@@ 1 Introduction
Large language models (LLMs) keep growing in width, depth, and sequence length [1, 2, 3, 4, 5, 6], pushing memory bandwidth and compute budgets to their limits in both training and deployment [7, 8, 9]. Quantization seeks to address these bottlenecks by storing weights and activations in a few bits and replacing many floating-point operations with much cheaper integer arithmetic [10]. There exist two mainstream workflows for quantization in the literature. Post-training quantization (PTQ) [11, 12, 13, 14, 15, 16] is attractive for its simplicity and zero retraining cost. However, at extremely low bit widths (e.g., 1–2 bits), it typically suffers from a severe accuracy drop. This failure often stems from compounding factors, including that calibration data are limited, per-tensor or per-group scaling is insufficient to realign codebooks, outlier channels dominate clipping ranges, etc. Quantization-aware training (QAT) [17, 18, 19, 20] inserts the quantizer into the training loop often with straight-through estimators [21], allowing the model to adapt to the quantized pattern. While current QAT methods outperform the PTQ methods under extremely low bit budgets, training a model from scratch with QAT is often prohibitively time-consuming and unstable since QAT methods still operate with tiny codebooks, offering insufficient freedom to match layer statistics.
@@
Recently, the iFairy framework [22] demonstrates the efficacy of transitioning from real-valued to complex-valued transformer architecture within the extremely low-bit quantization regime. By concurrently encoding magnitude and phase, complex weights introduce an additional degree of freedom that enhances modeling flexibility and representational capacity, thereby providing a theoretical basis for the superior performance of quantized complex-valued models. Crucially, iFairy adopts the fourth roots of unity on the unit circle ($\{\pm 1,\pm i\}$) as a compact codebook. This approach can fully exploit the 2 bit representation budget and hence achieve better alignment with weight statistics compared to real-valued counterparts, such as BitNet [23, 24, 25, 26], BitCPM4 [27], and ParetoQ [20], which typically rely on a ternary set $\{+1,0,-1\}$, resulting in the underutilization of the available 2-bit encoding space. Consequently, complex-valued Transformer models exhibit significantly improved accuracy retention at extremely low bit precision compared to state-of-the-art real-valued QAT approaches. Despite these benefits, iFairy models are trained from scratch with the QAT method. Reproducing such end-to-end training on larger models demands substantial time and compute, which hinders adoption when strong real-valued checkpoints already exist. In practice, many users require a path that reuses existing real-valued checkpoints and still obtains complex-domain benefits at an extremely low bit width.
@@
To address this problem, we propose Fairy2i, a universal framework for extremely-low bit QAT which converts a real-valued model into a complex-valued computation pattern and quantizes the converted model with efficient quantization paradigms in the complex domain. Specifically, we keep the pretrained real model, re-parameterize each real-valued linear layer in an equivalent widely-linear complex form, then continue training at extremely low bits in the QAT method with a simple phase-aware quantizer (e.g., PhaseQuant [22]), and further quantize the residual error to reduce the remaining gap at a tiny extra bit cost. The whole framework consists of three components, including widely-linear representation, phase-aware complex quantization, and recursive residual error quantization. In a widely-linear representation [28], we state that any real linear map on an even-dimensional space can be written exactly as the sum of a complex-linear part and a conjugate-linear part. We apply this to every linear layer of a real-valued model, generally including $\mathbf{Q}$ projection, $\mathbf{K}$ projection, $\mathbf{V}$ projection, and $\mathbf{O}$ projection in the self-attention block, as well as Up projection, Gate projection, and Down projection in the feed-forward network, taking LLaMA-alike architecture as an example. For attention, using the real part of the Hermitian inner product of queries and keys preserves the original scores. It can be proven that the forward computation is unchanged after the above transformation. On top of the widely-linear form, phase-aware complex quantization uses a very small complex codebook $\{\pm 1,\pm i\}$ with shared scaling factors across $\pm 1$ and $\pm i$ per group, respectively. We then train the transformed model with the QAT method while keeping full-precision master weights, so we can reuse existing real-valued checkpoints. To further reduce quantization error, we propose recursive residual error quantization, which also quantizes the residual error using the same phase-aware mechanism, recursively. The deployed weight is the sum of the extremely low bit representation of the original weight and its several successive quantization errors. This residual quantization trades a tiny extra code (one more 1–2 bits per weight, plus small per-group metadata) for a noticeable reduction in error.
@@ 2.1 Quantization for Large Language Models
Model quantization has become a de facto standard for compressing Large Language Models (LLMs). Existing methods can be broadly categorized into Post-Training Quantization (PTQ) [11] and Quantization-Aware Training (QAT).
@@
PTQ methods aim to compress models without extensive retraining, often using a small calibration dataset. Early works like AWQ [12] and GPTQ [11] focused on preserving the weights’ output activation distribution, achieving impressive results at 3-4 bits. More recent approaches like OmniQuant [29], QuIP [14], QuIP# [15] attempt to push the limit to 2 bits. However, PTQ methods typically suffer from severe performance degradation at extremely low bit-widths (e.g., $\leq 2$ bits) because the limited discrete codebook of real numbers (e.g., binary quantization of $\{+1,-1\}$ or ternary quantization of $\{+1,0,-1\}$) cannot adequately approximate the heavy-tailed distribution of LLM weights without adapting the network parameters themselves.
@@
To address the limitations of PTQ, QAT methods fine-tune the model parameters to adapt to the quantization noise. BitNet and its successor BitNet b1.58 introduced a 1-bit (binary $\{+1,-1\}$) and 1.58-bit (ternary $\{+1,0,-1\}$) architecture, respectively, demonstrating that LLMs can retain performance at extremely low precision if trained from scratch [30]. While effective, these methods often require training massive models from random initialization, incurring prohibitive computational costs. Although some works attempt to distill full-precision models into low-bit counterparts (e.g., LLM-QAT), the rigid topology of real-valued binary/ternary weights fundamentally limits the model’s capacity and optimization landscape, leading to instability or convergence issues when strictly enforced on pre-trained weights.
@@ 2.2 Complex-Valued Neural Networks
The integration of complex numbers into neural networks boasts a rich history, predominantly within domains where data possess intrinsic phase and magnitude, such as signal processing and imaging [31, 32, 33, 34, 35, 36]. By representing weights and activations in the complex domain, Complex-Valued Neural Networks (CVNNs) offer the theoretical potential to capture more intricate patterns and dependencies than their real-valued counterparts. However, despite this promise, the application of CVNNs to natural language processing—and to Large Language Models (LLMs) in particular—remains remarkably scarce.
@@
Recently, the iFairy [22] framework demonstrated the potential of complex-valued quantization for LLMs. By utilizing a phase-aware codebook $\{\pm 1,\pm i\}$ (the fourth roots of unity), iFairy showed that complex-valued Transformers could outperform real-valued binary networks at equivalent bit budgets. However, a major bottleneck of iFairy and similar methods is the lack of compatibility with existing real-valued pre-trained models like LLaMA [2, 3, 4], OPT [37], and Qwen [5]. They typically require training a complex-valued model from scratch, which is impractical for most practitioners given the scale of modern LLMs.
@@ 3.1 Widely-linear transformation from real domain to complex domain
where $(\cdot)^{\mathrm{H}}$ denotes the conjugate transpose. This formulation equals the original real dot-product scores when applied to stacked representations, i.e., $\operatorname{Re}(\mathbf{q}^{\mathrm{H}}\mathbf{k})=\operatorname{Re}(\mathbf{q})^{\top}\operatorname{Re}(\mathbf{k})+\operatorname{Im}(\mathbf{q})^{\top}\operatorname{Im}(\mathbf{k})=\tilde{\mathbf{q}}^{\top}\tilde{\mathbf{k}}$. The softmax and value aggregation act separately on the real and imaginary parts, as the attention score is a real number; the output projection uses the same widely-linear mapping. Note that the self-attention in a widely-linear form is compatible with highly-optimized FlashAttention kernels [38, 39, 40] since we can unstack the complex-valued $\mathbf{q},\mathbf{k},\mathbf{v}$ into a longer real-valued vector, respectively.
@@ 3.3 Recursive residual quantization
[width=0.7]Figure/Fairy2i-recursive
@@
We maintain full-precision masters for $W$ and compute $\sum_{t}\widehat{W}^{(t)}$ in the forward pass. The projection $b(\cdot)$ uses the hard nearest-codeword rule; in the backward pass, we apply a straight-through estimator so that gradients flow to the masters. Scales are recomputed per step (or updated as moving averages) and are differentiable through their dependence on the residuals. We apply the same procedure independently to $\mathbf{U}$ and $\mathbf{W}$ in every widely-linear layer.
@@ 3.4 Storage Efficiency and Parallel Multiplication-Free Inference
Beyond multiplication-free arithmetic, the discreteness of our 2-bit weights enables highly efficient inference via Look-Up Tables (LUTs), particularly on CPUs [41]. Following the principles of optimized kernels like BitNet.cpp [42] and T-MAC [43], we can pack groups of four 2-bit complex weights into a single 8-bit index. By pre-computing the partial outcomes of these weight combinations with INT8 activations, the inner loop of matrix multiplication reduces to a simple table fetch and accumulation. This approach significantly accelerates execution by minimizing arithmetic instructions while retaining mathematical exactness.
@@ 4.1 Experimental Setup
    • Real-Binary (1-bit): Weights are quantized to $\{+1,-1\}$, similar to BitNet [23].
    • Real-Ternary (1.58-bit): Weights are quantized to $\{+1,0,-1\}$, similar to BitNet b1.58 [26].
  • PTQ Baselines: We also include results from GPTQ [11], QuIP# [15] and AQLM [44] for reference.
@@
We evaluate the quantized models on both language modeling capabilities and downstream reasoning tasks using the lm-eval-harness framework [45].
  • Perplexity (PPL): We report perplexity on the validation set of C4 [46].
  • Zero-shot Tasks: We evaluate zero-shot accuracy on several common sense reasoning benchmarks: ARC-Easy [47], ARC-Challenge [47], HellaSwag [48], PIQA [49], and Winogrande [50].
@@
We implement Fairy2i using PyTorch. The real-valued pretrained weights are first transformed into the widely-linear complex form. We then continue pre-training with the QAT method on a subset of the RedPajama dataset [51] for 30 billion tokens. We use the AdamW [52] optimizer with a Warmup-Stable-Decay (WSD) learning rate schedule [53, 27]. The global batch size is set to 1 million tokens. Unless otherwise specified, we report results for two configurations of Fairy2i:
@@ 4.2 Main Results
Table 1: C4 Perplexity and Zero-shot Accuracy comparison on LLaMA-2 7B. Avg. denotes the average accuracy across the 5 tasks. Results for AQLM, QuIP#, and GPTQ are sourced from [44].
@@
GPTQ [11] 3 10.61 58.46 31.06 45.21 71.49 59.19 53.08
QuIP# [15] 2 11.01 55.56 28.84 42.94 71.38 62.43 52.23
AQLM [44] 2 8.54 63.68 32.76 49.55 74.76 65.67 57.28

### v3

@@ Abstract
Large language models (LLMs) have revolutionized artificial intelligence, yet their massive memory and computational demands necessitate aggressive quantization, increasingly pushing representations toward the theoretical limit of a single bit. While complex-valued LLMs, such as iFairy, offer a superior chance for low-bit representation compared to real-valued counterparts, they require training from scratch, preventing the utilization of the vast ecosystem of pre-trained real-valued foundation models. Here we present Fairy2i, a universal framework that transforms pre-trained real-valued layers into an equivalent widely-linear complex form, enabling extremely low-bit quantization while reusing existing checkpoints. By proving a lossless mathematical equivalence between real and widely-linear maps, we convert standard Transformers into the complex domain and employ a phase-aware quantization scheme with a highly efficient codebook of fourth roots of unity ($\{\pm 1,\pm i\}$). Furthermore, we introduce a recursive residual quantization mechanism that iteratively minimizes quantization error, allowing inference to proceed via efficient multiplication-free accumulation. We demonstrate that Fairy2i restores the performance of LLaMA-2 7B at an effective 2-bit precision to levels nearly comparable with full-precision baselines, significantly outperforming state-of-the-art real-valued binary and ternary quantization methods. This work bridges the gap between the representational efficiency of complex-valued arithmetic and the practical utility of pre-trained models, paving a new way for efficient inference on commodity hardware. We open-source the Fairy2i model and code at https://huggingface.co/PKU-DS-LAB/Fairy2i-W2 and https://github.com/PKULab1806/Fairy2i-W2.
@@ 1 Introduction
Large language models (LLMs) keep growing in width, depth, and sequence length [undef, undefa, undefb, undefc, undefd, undefe], pushing memory bandwidth and compute budgets to their limits in both training and deployment [undeff, undefg, undefh]. Quantization seeks to address these bottlenecks by storing weights and activations in a few bits and replacing many floating-point operations with much cheaper integer arithmetic [undefi]. There exist two mainstream workflows for quantization in the literature. Post-training quantization (PTQ) [undefj, undefk, undefl, undefm, undefn, undefo] is attractive for its simplicity and zero retraining cost. However, at extremely low bit widths (e.g., 1–2 bits), it typically suffers from a severe accuracy drop. This failure often stems from compounding factors, including that calibration data are limited, per-tensor or per-group scaling is insufficient to realign codebooks, outlier channels dominate clipping ranges, etc. Quantization-aware training (QAT) [undefp, undefq, undefr, undefs] inserts the quantizer into the training loop often with straight-through estimators [undeft], allowing the model to adapt to the quantized pattern. While current QAT methods outperform the PTQ methods under extremely low bit budgets, training a model from scratch with QAT is often prohibitively time-consuming and unstable since QAT methods still operate with tiny codebooks, offering insufficient freedom to match layer statistics.
@@
Recently, the iFairy framework [undefu] demonstrates the efficacy of transitioning from real-valued to complex-valued transformer architecture within the extremely low-bit quantization regime. By concurrently encoding magnitude and phase, complex weights introduce an additional degree of freedom that enhances modeling flexibility and representational capacity, thereby providing a theoretical basis for the superior performance of quantized complex-valued models. Crucially, iFairy adopts the fourth roots of unity on the unit circle ($\{\pm 1,\pm i\}$) as a compact codebook. This approach can fully exploit the 2 bit representation budget and hence achieve better alignment with weight statistics compared to real-valued counterparts, such as BitNet [undefv, undefw, undefx, undefy], BitCPM4 [undefz], and ParetoQ [undefs], which typically rely on a ternary set $\{+1,0,-1\}$, resulting in the underutilization of the available 2-bit encoding space. Consequently, complex-valued Transformer models exhibit significantly improved accuracy retention at extremely low bit precision compared to state-of-the-art real-valued QAT approaches. Despite these benefits, iFairy models are trained from scratch with the QAT method. Reproducing such end-to-end training on larger models demands substantial time and compute, which hinders adoption when strong real-valued checkpoints already exist. In practice, many users require a path that reuses existing real-valued checkpoints and still obtains complex-domain benefits at an extremely low bit width.
@@
To address this problem, we propose Fairy2i, a universal framework for extremely-low bit QAT which converts a real-valued model into a complex-valued computation pattern and quantizes the converted model with efficient quantization paradigms in the complex domain. Specifically, we keep the pretrained real model, re-parameterize each real-valued linear layer in an equivalent widely-linear complex form, then continue training at extremely low bits in the QAT method with a simple phase-aware quantizer (e.g., PhaseQuant [undefu]), and further quantize the residual error to reduce the remaining gap at a tiny extra bit cost. The whole framework consists of three components, including widely-linear representation, phase-aware complex quantization, and recursive residual error quantization. In a widely-linear representation [undefaa], we state that any real linear map on an even-dimensional space can be written exactly as the sum of a complex-linear part and a conjugate-linear part. We apply this to every linear layer of a real-valued model, generally including $\mathbf{Q}$ projection, $\mathbf{K}$ projection, $\mathbf{V}$ projection, and $\mathbf{O}$ projection in the self-attention block, as well as Up projection, Gate projection, and Down projection in the feed-forward network, taking LLaMA-alike architecture as an example. For attention, using the real part of the Hermitian inner product of queries and keys preserves the original scores. It can be proven that the forward computation is unchanged after the above transformation. On top of the widely-linear form, phase-aware complex quantization uses a very small complex codebook $\{\pm 1,\pm i\}$ with shared scaling factors across $\pm 1$ and $\pm i$ per group, respectively. We then train the transformed model with the QAT method while keeping full-precision master weights, so we can reuse existing real-valued checkpoints. To further reduce quantization error, we propose recursive residual error quantization, which also quantizes the residual error using the same phase-aware mechanism, recursively. The deployed weight is the sum of the extremely low bit representation of the original weight and its several successive quantization errors. This residual quantization trades a tiny extra code (one more 1–2 bits per weight, plus small per-group metadata) for a noticeable reduction in error.
@@ 2.1 Quantization for Large Language Models
Model quantization has become a de facto standard for compressing Large Language Models (LLMs). Existing methods can be broadly categorized into Post-Training Quantization (PTQ) [undefj] and Quantization-Aware Training (QAT).
@@
PTQ methods aim to compress models without extensive retraining, often using a small calibration dataset. Early works like AWQ [undefk] and GPTQ [undefj] focused on preserving the weights’ output activation distribution, achieving impressive results at 3-4 bits. More recent approaches like OmniQuant [undefab], QuIP [undefm], QuIP# [undefn] attempt to push the limit to 2 bits. However, PTQ methods typically suffer from severe performance degradation at extremely low bit-widths (e.g., $\leq 2$ bits) because the limited discrete codebook of real numbers (e.g., binary quantization of $\{+1,-1\}$ or ternary quantization of $\{+1,0,-1\}$) cannot adequately approximate the heavy-tailed distribution of LLM weights without adapting the network parameters themselves.
@@
To address the limitations of PTQ, QAT methods fine-tune the model parameters to adapt to the quantization noise. BitNet and its successor BitNet b1.58 introduced a 1-bit (binary $\{+1,-1\}$) and 1.58-bit (ternary $\{+1,0,-1\}$) architecture, respectively, demonstrating that LLMs can retain performance at extremely low precision if trained from scratch [undefac]. While effective, these methods often require training massive models from random initialization, incurring prohibitive computational costs. Although some works attempt to distill full-precision models into low-bit counterparts (e.g., LLM-QAT), the rigid topology of real-valued binary/ternary weights fundamentally limits the model’s capacity and optimization landscape, leading to instability or convergence issues when strictly enforced on pre-trained weights.
@@ 2.2 Complex-Valued Neural Networks
The integration of complex numbers into neural networks boasts a rich history, predominantly within domains where data possess intrinsic phase and magnitude, such as signal processing and imaging [undefad, undefae, undefaf, undefag, undefah, undefai]. By representing weights and activations in the complex domain, Complex-Valued Neural Networks (CVNNs) offer the theoretical potential to capture more intricate patterns and dependencies than their real-valued counterparts. However, despite this promise, the application of CVNNs to natural language processing—and to Large Language Models (LLMs) in particular—remains remarkably scarce.
@@
Recently, the iFairy [undefu] framework demonstrated the potential of complex-valued quantization for LLMs. By utilizing a phase-aware codebook $\{\pm 1,\pm i\}$ (the fourth roots of unity), iFairy showed that complex-valued Transformers could outperform real-valued binary networks at equivalent bit budgets. However, a major bottleneck of iFairy and similar methods is the lack of compatibility with existing real-valued pre-trained models like LLaMA [undefa, undefb, undefc], OPT [undefaj], and Qwen [undefd]. They typically require training a complex-valued model from scratch, which is impractical for most practitioners given the scale of modern LLMs.
@@ 3.1 Widely-linear transformation from real domain to complex domain
where $(\cdot)^{\mathrm{H}}$ denotes the conjugate transpose. This formulation equals the original real dot-product scores when applied to stacked representations, i.e., $\operatorname{Re}(\mathbf{q}^{\mathrm{H}}\mathbf{k})=\operatorname{Re}(\mathbf{q})^{\top}\operatorname{Re}(\mathbf{k})+\operatorname{Im}(\mathbf{q})^{\top}\operatorname{Im}(\mathbf{k})=\tilde{\mathbf{q}}^{\top}\tilde{\mathbf{k}}$. The softmax and value aggregation act separately on the real and imaginary parts, as the attention score is a real number; the output projection uses the same widely-linear mapping. Note that the self-attention in a widely-linear form is compatible with highly-optimized FlashAttention kernels [undefak, undefal, undefam] since we can unstack the complex-valued $\mathbf{q},\mathbf{k},\mathbf{v}$ into a longer real-valued vector, respectively.
@@ 3.3 Recursive residual quantization
[width=0.7]Figure/recursive-darker
@@
We maintain full-precision masters for $W$ and compute $\sum_{t}\widehat{W}^{(t)}$ in the forward pass. The projection $b(\cdot)$ uses the hard nearest-codeword rule; in the backward pass, we apply a straight-through estimator so that gradients flow to the masters. Scales are recomputed per step and are differentiable through their dependence on the residuals. We apply the same procedure independently to $\mathbf{U}$ and $\mathbf{W}$ in every widely-linear layer.
@@ 3.4 Storage Efficiency and Parallel Multiplication-Free Inference
Beyond multiplication-free arithmetic, the discreteness of our 2-bit weights enables highly efficient inference via Look-Up Tables (LUTs), particularly on CPUs [undefan]. Following the principles of optimized kernels like BitNet.cpp [undefao] and T-MAC [undefap], we can pack groups of four 2-bit complex weights into a single 8-bit index. By pre-computing the partial outcomes of these weight combinations with INT8 activations, the inner loop of matrix multiplication reduces to a simple table fetch and accumulation. This approach significantly accelerates execution by minimizing arithmetic instructions while retaining mathematical exactness.
@@ 4.1 Experimental Setup
    • Real-Binary (1-bit): Weights are quantized to $\{+1,-1\}$, similar to BitNet [undefv].
    • Real-Ternary (1.58-bit): Weights are quantized to $\{+1,0,-1\}$, similar to BitNet b1.58 [undefy].
  • PTQ Baselines: We also include results from GPTQ [undefj], QuIP# [undefn] and AQLM [undefaq] for reference.
@@
We evaluate the quantized models on both language modeling capabilities and downstream reasoning tasks using the lm-eval-harness framework [undefar].
  • Perplexity (PPL): We report perplexity on the validation set of C4 [undefas].
  • Zero-shot Tasks: We evaluate zero-shot accuracy on several common sense reasoning benchmarks: ARC-Easy [undefat], ARC-Challenge [undefat], HellaSwag [undefau], PIQA [undefav], and Winogrande [undefaw].
@@
We implement Fairy2i using PyTorch. The real-valued pretrained weights are first transformed into the widely-linear complex form. We then continue pre-training with the QAT method on a subset of the RedPajama dataset [undefax] for 30 billion tokens. We use the AdamW [undefay] optimizer with a Warmup-Stable-Decay (WSD) learning rate schedule [undefaz, undefz]. The global batch size is set to 1 million tokens. Unless otherwise specified, we report results for two configurations of Fairy2i:
@@ 4.2 Main Results
Table 1: C4 Perplexity and Zero-shot Accuracy comparison on LLaMA-2 7B. Avg. denotes the average accuracy across the 5 tasks. Results for AQLM, QuIP#, and GPTQ are sourced from [undefaq].
@@
GPTQ [undefj] 3 10.61 58.46 31.06 45.21 71.49 59.19 53.08
QuIP# [undefn] 2 11.01 55.56 28.84 42.94 71.38 62.43 52.23
AQLM [undefaq] 2 8.54 63.68 32.76 49.55 74.76 65.67 57.28